Please solve this question, and show steps. Thank you so much.

A propeller is modeled as five identical uniform rods extending radially from its axis. The length and mass of each rod are 0​​​​​​​
Answer #1

Given) m=5, 2=0.809m, m=2.7 L N-599 rpm Sol.) rotational KE = n I zw? → Kern mecan) + XE = 2 ( 2.75) 0.8.92 ( 8T5 gg) 2 KE= 9